# The SABRE South Collaboration

### v1

@@ Abstract
Ongoing efforts are underway from a number of experimental collaborations aiming to unambiguously confirm or refute the DAMA signal in experiments using the same target material [9, 10, 3]. Such a determination is of great importance; to date, no satisfactory explanation for the signal consistent with other experimental results has been given. Recent works have aimed to highlight that aspects of the modulation reported by DAMA could be explained by a particular analysis technique that they apply to their data [2, 11, 12]. Here, we show that a more detailed exploration of this effect reveals strong inconsistencies between the signal-like component ‘induced’ by the analysis technique and the results reported by DAMA. In particular, we highlight the incorrect phase of the induced signal, the substantial reduction in amplitude coming from a re-assessment of their likely backgrounds, and the non-constant nature of the signal amplitude in time. We additionally explore an alternative statistical treatment of the data used by DAMA, showing that whilst this also has the potential to produce a bias under their assumptions on their background model, the bias produced is once again inconsistent with the reported results. Finally, we examine the interplay between the DAMA analysis technique and a genuinely modulating component, be it from a dark matter signal or a source of background.
@@
We find that the induced modulation effect comes from the application of the DAMA analysis method in the presence of a background rate that is not constant in time. As identified in Ref. [2], the reported rate of single-hit events in the DAMA experiment was found to decrease between phase 1 and phase 2 of the experiment [^1] Phase 1 ran from September 2003 to August 2010, whilst phase 2 ran from November 2011 to October 2019, with PMT upgrades occuring in between. [14, 15]. Background modelling from the NaI(Tl) experiments COSINE-100 [16] and ANAIS-112 [17], as well as background simulation from the SABRE South experiment [13], suggests that time variation in the 2-6 keV DAMA region of interest (ROI) over the course of the experiment [^2] Whilst other cosmogenically activated isotopes do exhibit time variation, these timescales are either much longer or much shorter than the experimental runtime. is expected from the decay of the radioisotopes 3H (tritium, half-life 12.3 years) and 210Pb (half-life 22.3 years). As such, we scale the initial activities of these background model components from the values used for the SABRE South simulation to match the upper limits reported on each by DAMA [18]. These are 0.09 and 0.03 mBq/kg for 3H and 210Pb, respectively. In addition, we include a component of 210Pb within the Teflon wrap of the crystals. This is a conservative choice given some evidence from the characterisation of the SABRE North crystal, denoted NaI-33, that this activity could be higher than previously assumed [19, 13], although we note that its inclusion has little impact. The upper limit on the 210Pb activity from the SABRE North measurement is used to scale this particular surface component in the SABRE South simulation, correcting for the different crystal surface area-to-volume ratios. All other initial activities are set to the values reported in the SABRE South simulation paper [13]. Given that the induced modulation effect comes from the decaying components, we consider this assumption to be reasonable. Our assumption is further supported by the expected similarity in crystal purity and composition between the DAMA crystals and the SABRE South crystals [20, 21].
@@
Figure 1: Top: a single realisation of a toy DAMA/LIBRA dataset, generated as described in the text over the entire experimental exposure. The exposure-normalised number of counts in 10 bins across each approximately annual cycle – depicted with dashed vertical lines – is plotted in red. Black horizontal lines denote the average of the normalised rate over each annual cycle. Time bins/annual cycles are plotted such that they appear continuous/contiguous (see text). Bottom: residuals distribution for the above dataset, calculated for each bin as described in the text. The red line shows the result of fitting an oscillating function to the residuals, as described in the text.
@@
where $\langle R_{i}\rangle_{j}$ denotes the event rate averaged over all bins of the $j$th cycle. We plot these residuals for the same toy dataset in Fig. 1. The DAMA methodology then fits a modulating function of the form $\mathcal{A}\cos\left[\frac{2\pi}{T}t_{i}+\phi\right]$ to the ${R}_{ij}$, with $t_{i}$ being the time at the centre of the $i$th time bin and $\mathcal{A}$, $\phi$ and $T$ being the amplitude, phase and period of the modulation, respectively. The motivation behind this approach is to provide a measure of the modulating event rate around a constant background within each annual cycle. However, as is evident from Fig. 1, in the presence of a time-varying background this method produces a bias. This is the induced modulation effect; where the background rate is monotonically changing with time, such a procedure will produce a sawtooth-like set of residuals. The period of these residuals is determined by the length of the averaging period, the amplitude by the magnitude of the variation of the rate around the average within a cycle, and the phase by the nature of the time variation with respect to the average. We show the result of performing this fit to this particular set of residuals in Fig. 1.
@@
Following a similar approach to Ref. [2], we study the expected distribution of amplitudes, phases, and periods under the application of this analysis procedure over a large number (1000) of toy realisations of the DAMA dataset. Each dataset is generated over the 15 annual cycles across DAMA/LIBRA phases 1 and 2, with the ROI being restricted to 2-6 keV for comparison with the combined residuals analysis presented by DAMA [1]. Above the energy threshold of 2 keV 5551 keV for DAMA/LIBRA phase 2., DAMA state that they have full rejection of background events from PMT noise [20], and as such we do not consider this component here. For each toy, a fit of the form $\mathcal{A}\cos\left[\frac{2\pi}{T}t_{i}+\phi\right]$ is performed as a $\chi^{2}$ minimisation using the residuals ${R}_{ij}$ and their uncertainties. We show the resultant distributions of the three fit parameters in orange in Fig. 2. Similarly to Ref. [2], we find that the amplitudes determined from the fits are clustered around the amplitude reported by DAMA, whilst the periods are clustered around 12.1 months, just above the fit value of 12 months reported by DAMA [1]. However, crucially, we find that the phases determined from the fits – relative to the expected dark matter modulation phase $\phi_{0}$ corresponding to June 2nd – are clustered close to $\pi$, i.e. they are close to being in antiphase with the modulation signal that would originate from a galactic dark matter source. A similar effect was observed in Ref. [2], and we emphasise here its importance: such an induced modulation effect cannot explain the result of the fit presented by DAMA. Given the fixed locations of the annual cycles used for the averaging procedure, an induced modulation effect giving an approximately consistent phase with the DAMA fit could only come from a monotonically increasing background rate, which seems unphyiscal.
@@
We now study the impact of this lower 3H background on the induced modulation effect, by performing the same series of toy fits as before, this time setting the initial 3H activity in the toy background model to 0.031 mBq/kg. We show the resultant distributions of $\mathcal{A}$, $\phi$ and $T$ in Fig. 2, this time in blue. Whilst the distributions of both the period and the phase determined in the fits are similar to before, we find the distribution of amplitudes to be shifted to much lower values, by a factor $\sim$2. This reflects the relative rate reduction across each annual cycle being lower as a consequence of the initial 3H activity being lower. We view this as further evidence against the hypothesis that the DAMA signal is an artefact of their analysis technique; with what we see as a more likely, but still over-attributed estimation of their 3H background, neither the amplitude nor the phase of the induced modulation are consistent with their reported results.
@@
Here, $\mathcal{E}_{i}$ is the exposure of the $i$th time bin, $\Delta E_{j}$ is the width of the $j$th energy bin and $\epsilon_{j}$ is the energy-dependent efficiency at the centre of that energy bin. The time at the centre of the $i$th time bin is given by $t_{i}$, $T$ is the period of the oscillating component – fixed here to 12 months – and $t_{0}$ is the time of maximum modulation amplitude, fixed here to 152.5 days. The free parameters of the fit are $c_{j}$ and $m_{j}$, controlling the amplitude of a constant and modulating component for each energy bin.
@@
It is interesting to observe where the energy dependence of this fit in the presence of such high concentrations of these isotopes comes from. Given that there is no notion of yearly-averaged background subtraction present in this fitting methodology, it would be misleading to attribute what we are seeing to the same induced modulation effect seen with the residuals. To understand why an energy dependence is observed, it is helpful to visualise the distribution of $N_{ij}$ and $\mu_{ij}$. Figure 4 shows normalised $N_{ij}$ and $\mu_{ij}$, evaluated at the best-fit values $\hat{c}_{j}$ and $\hat{m}_{j}$, for both the first and last energy bins, for a particular realisation of a toy dataset using the DAMA-reported upper limit on 3H activity. As can be seen, a strong decay in counts across time bins is seen for the first energy bin, with an almost imperceptible level of decay for the last energy bin. We attribute this to the fact that the energy spectra of the components with the strongest reduction in count rates across the live time of the experiment – 3H and 210Pb – are concentrated in the lowest-energy bins. In neither case is a modulation seen. We see that $\hat{m}_{j}$ taking a non-zero and negative value in the first energy bin is simply a consequence of this value minimising the deviation between observed and expected counts across bins.
Figure 5: The result of performing the same fit to a single realisation of a toy DAMA/LIBRA dataset as in Fig. 1, but this time partitioning the data into 5 equal time periods across the exposure and performing the fit separately for each. As can be seen, the amplitude of the induced modulation decays over time.
An especially curious feature of the DAMA/LIBRA modulation signal is its persistence over time. Such behaviour is cited by DAMA when they argue that the modulation of their residual rate – seen to be constant across the entirety of phases 1 and 2 of the experiment [1] – could be due to the scattering of galactic dark matter. Clearly, though, the induced modulation effect coming from the presence of an exponentially falling background rate produces a modulation-like signature that is not constant in time. This is made evident in Fig. 5, where the same fit is performed as in Fig. 1, this time fitting within 3 slices of time across the full exposure. It is clear that the amplitude of the fit is lower for later slices. It seems difficult, therefore, to attribute the DAMA signal to an induced modulation coming from exponentially decaying background activity.
Figure 6: The result of performing the same fit to a single realisation of a toy DAMA/LIBRA dataset as in Fig. 1, but this time including a component with a genuinely oscillating event rate, as described in the text. We compare a fit to only the modulating component of the datset (blue), versus the full dataset (orange), under two assumptions on the initial activities of 3H and 210Pb. Top: initial activities of 3H and 210Pb are set to those reported by DAMA. Bottom: initial activities of 3H and 210Pb are set to zero.

### v2

@@ Abstract
Ongoing efforts are underway from a number of experimental collaborations aiming to unambiguously confirm or refute the DAMA signal in experiments using the same target material, including already published initial results from the COSINE-100 [12] and ANAIS-112 [13] experiments. Such a determination is of great importance; to date, no satisfactory explanation for the signal consistent with other experimental results has been given. Recent works have debated whether aspects of the modulation reported by DAMA could be explained by a particular analysis technique that they apply to their data [2, 14]. Here, we show that a more detailed exploration of this effect reveals strong inconsistencies between the signal-like component ‘induced’ by the analysis technique and the results reported by DAMA. In particular, we highlight the incorrect phase of the induced signal, the substantial reduction in amplitude coming from a re-assessment of their likely backgrounds, and the non-constant nature of the signal amplitude in time. We additionally explore an alternative statistical treatment of the data used by DAMA, showing that whilst this also has the potential to produce a bias under their assumptions on their background model, the bias produced is once again inconsistent with the reported results. Finally, we examine the interplay between the DAMA analysis technique and a genuinely modulating component, be it from a dark matter signal or a source of background.
@@
We find that the induced modulation effect comes from the application of the DAMA analysis method in the presence of a background rate that is not constant in time. As identified in Ref. [2], the reported rate of single-hit events in the DAMA experiment was found to decrease between phase 1 and phase 2 of the experiment [^2] Phase 1 ran from September 2003 to August 2010, whilst phase 2 ran from November 2011 to October 2019, with PMT upgrades occuring in between. [16, 17]. Background modelling from the NaI(Tl) experiments COSINE-100 [18] and ANAIS-112 [19], as well as background simulation from the SABRE South experiment [15], suggests that time variation in the 2-6 keV DAMA region of interest (ROI) over the course of the experiment [^3] Whilst other cosmogenically activated isotopes do exhibit time variation, these timescales are either much longer or much shorter than the experimental runtime. is expected from the decay of the radioisotopes 3H (tritium, half-life 12.3 years) and 210Pb (half-life 22.3 years). As such, we scale the initial activities of these background model components from the values used for the SABRE South simulation to match the upper limits reported on each by DAMA [20]. These are 0.09 and 0.03 mBq/kg for 3H and 210Pb, respectively. In addition, we include a component of 210Pb within the Teflon wrap of the crystals. This is a conservative choice given some evidence from the characterisation of the SABRE North crystal, denoted NaI-33, that this activity could be higher than previously assumed [21, 15], although we note that its inclusion has little impact. The upper limit on the 210Pb activity from the SABRE North measurement is used to scale this particular surface component in the SABRE South simulation, correcting for the different crystal surface area-to-volume ratios. All other initial activities are set to the values reported in the SABRE South simulation paper [15]. Given that the induced modulation effect comes from the decaying components, we consider this assumption to be reasonable. Our assumption is further supported by the expected similarity in crystal purity and composition between the DAMA crystals and the SABRE South crystals [22, 23]. We do not consider the effect of muon-induced neutron backgrounds in this work [24]. .
@@
Figure 1: Top: a single realisation of a toy DAMA/LIBRA dataset, generated as described in the text over the entire experimental exposure. The exposure-normalised number of counts in 10 bins across each approximately annual cycle – depicted with dashed vertical lines – is plotted in red. Black horizontal lines denote the average of the normalised rate over each annual cycle. Time bins/annual cycles are plotted such that they appear continuous/contiguous (see text). Bottom: residuals distribution for the above dataset, calculated for each bin as described in the text. The red line shows the result of fitting an oscillating function to the residuals, as described in the text. Error bars on the residuals are calculated by combining the Poisson errors on the counts in each bin with the error on the mean counts subtracted within each annual cycle.
@@
where $\langle R_{i}\rangle_{j}$ denotes the event rate averaged over all bins of the $j$th cycle. We plot these residuals for the same toy dataset in Fig. 1. The DAMA methodology [^6] DAMA also performs fits with a fixed period and phase; throughout this work, we employ a fit where all three parameters freely float. then fits a modulating function of the form $\mathcal{A}\cos\left[\frac{2\pi}{T}t_{i}+\phi\right]$ to the ${R}_{ij}$, with $t_{i}$ being the time at the centre of the $i$th time bin and $\mathcal{A}$, $\phi$ and $T$ being the amplitude, phase and period of the modulation, respectively. The motivation behind this approach is to provide a measure of the modulating event rate around a constant background within each annual cycle. However, as is evident from Fig. 1, in the presence of a time-varying background this method produces a bias. This is the induced modulation effect; where the background rate is monotonically changing with time, such a procedure will produce a sawtooth-like set of residuals. The period of these residuals is determined by the length of the averaging period, the amplitude by the magnitude of the variation of the rate around the average within a cycle, and the phase by the nature of the time variation with respect to the average. We show the result of performing this fit to this particular set of residuals in Fig. 1.
@@
Following a similar approach to Ref. [2], we study the expected distribution of amplitudes, phases, and periods under the application of this analysis procedure over a large number (1000) of toy realisations of the DAMA dataset. Each dataset is generated over the 15 annual cycles across DAMA/LIBRA phases 1 and 2, with the ROI being restricted to 2-6 keV for comparison with the combined residuals analysis presented by DAMA [1]. Above the energy threshold of 2 keV7771 keV for DAMA/LIBRA phase 2., DAMA state that they have full rejection of background events from PMT noise [22], and as such we do not consider this component here. For each toy, a fit of the form $\mathcal{A}\cos\left[\frac{2\pi}{T}t_{i}+\phi\right]$ is performed as a $\chi^{2}$ minimisation using the residuals ${R}_{ij}$ and their uncertainties. We show the resultant distributions of the three fit parameters in orange in Fig. 2. Similarly to Ref. [2], we find that the amplitudes $\mathcal{A}$ determined from the fits are clustered around the amplitude reported by DAMA, whilst the periods $T$ are clustered around 12.1 months, just above the fit value of 12 months reported by DAMA [1]. However, crucially, we find that the phases $\phi$ determined from the fits – relative to the expected dark matter modulation phase $\phi_{0}$ corresponding to June 2nd – are clustered close to $\pi$, i.e. they are close to being in antiphase with the modulation signal that would originate from a galactic dark matter source. A similar effect was observed in Ref. [2], and we emphasise here its importance: such an induced modulation effect cannot explain the result of the fit presented by DAMA. Given the fixed locations of the annual cycles used for the averaging procedure, an induced modulation effect giving an approximately consistent phase with the DAMA fit could only come from a monotonically increasing background rate, which seems unphyiscal.
@@
We now study the impact of this lower 3H background on the induced modulation effect, by performing the same series of toy fits as before, this time setting the initial 3H activity in the toy background model to 0.031 mBq/kg. We show the resultant distributions of $\mathcal{A}$, $\phi$ and $T$ in Fig. 2, this time in blue. Whilst the distributions of both the period and the phase determined in the fits are similar to before, we find the distribution of amplitudes to be shifted to much lower values, by a factor $\sim$2. This reflects the relative rate reduction across each annual cycle being lower as a consequence of the initial 3H activity being lower. We view this as further evidence against the hypothesis that the DAMA signal is an artefact of their determination of the modulation; with what we see as a more likely, but still over-attributed estimation of their 3H background, neither the amplitude nor the phase of the induced modulation are consistent with their reported results. We note that both 3H activity assumptions lead to a poor goodness of fit metric: the mean $\chi^{2}/d.o.f.$ values over the 1000 toys are found to be 471 / 147 and 413 / 147 for the DAMA-reported and revised activities, respectively. This can be compared with 130 / 155 for DAMA’s fit [1] [^8] It should be noted that this is for a fit with a fixed period and phase, and also includes data from DAMA/NaI.. The fact that a better fit is found for the revised 3H activity case can be attributed to the fact that there is a less dramatic decrease in modulation amplitude in this case.
@@
Here, $\mathcal{E}_{i}$ is the exposure of the $i$th time bin, $\Delta E_{j}$ is the width of the $j$th energy bin and $\epsilon_{j}$ is the energy-dependent efficiency at the centre of that energy bin. The time at the centre of the $i$th time bin is given by $t_{i}$, $T$ is the period of the oscillating component – fixed here to 365.25 days – and $t_{0}$ is the time of maximum modulation amplitude, fixed here to 152.5 days. The free parameters of the fit are $c_{j}$ and $m_{j}$, controlling the amplitude of a constant and modulating component for each energy bin.
@@
It is interesting to observe where the energy dependence of this fit in the presence of such high concentrations of these isotopes comes from. Given that there is no notion of yearly-averaged background subtraction present in this fitting methodology, it would be misleading to attribute what we are seeing to the same induced modulation effect seen with the residuals. To understand why an energy dependence is observed, it is helpful to visualise the distribution of the obeserved counts $N_{ij}$ and the expected counts $\mu_{ij}$. Figure 4 shows normalised $N_{ij}$ and $\mu_{ij}$, evaluated at the best-fit values $\hat{c}_{j}$ and $\hat{m}_{j}$, for both the first and last energy bins, for a particular realisation of a toy dataset using the DAMA-reported upper limit on 3H activity. As can be seen, a strong decay in counts across time bins is seen for the first energy bin, with an almost imperceptible level of decay for the last energy bin. We attribute this to the fact that the energy spectra of the components with the strongest reduction in count rates across the live time of the experiment – 3H and 210Pb – are concentrated in the lowest-energy bins. In neither case is a modulation seen. We see that $\hat{m}_{j}$ taking a non-zero and negative value in the first energy bin is simply a consequence of this value minimising the deviation between observed and expected counts across bins.
@@
We note that DAMA considered including a non-oscillatory time dependence in this approach by including in the square brackets of Eqn. 5 a term $l_{j}t_{i}$, i.e. a linear decrease (or increase) in rate over time in each energy bin, finding little change in their fit results [22]. We repeat the analysis procedure used to produce Fig. 3, this time including this linear component in our fit function for $\mu_{ij}$. We show the result of this in Fig. 5, where it is clear that we still see a strong inconsistency between our toy fits and the reported DAMA results. We find that whilst the sign of $\hat{m}_{j}$ in the lower energy bins is still opposite to that reported in the DAMA fits, the amplitude is much smaller compared to our toy fits without the linear component, and we see much less difference in our fit results for the different 3H activity assumptions. This can be attributed to the linear component in each energy bin allowing for the deviation between observed and expected counts to be minimised with a much smaller fitted $\hat{m}_{j}$, given that $\hat{l}_{j}$ has the freedom to fit out differently for the different 3H activity cases. We conclude that this fitting methodology, in the presence of time-varying 210Pb and 3H backgrounds, cannot produce results consistent with those reported by DAMA, implying a feature in their data not captured by our toy model.
Figure 5: The result of performing the same analysis as in Fig. 3, this time including a linear term in Eqn. 5. We once again compare with the DAMA-reported results [1], unchanged with the inclusion of the linear term [22].
Figure 6: The result of performing the same fit as in Fig. 1 to a single realisation of a toy DAMA/LIBRA dataset, this time partitioning the data into 3 equal time periods and performing the fit separately for each. As can be seen, the amplitude of the induced modulation decays over time. We compare with a fit to the dataset without any partitioning (dashed black).
An especially curious feature of the DAMA/LIBRA modulation signal is its persistence over time. Such behaviour is cited by DAMA when they argue that the modulation of their residual rate – seen to be constant across the entirety of phases 1 and 2 of the experiment [1] – could be due to the scattering of galactic dark matter. Clearly, though, the induced modulation effect coming from the presence of an exponentially falling background rate produces a modulation-like signature that is not constant in time. This is made evident in Fig. 6, where the same fitting procedure is performed as in Fig. 1, this time fitting within 3 slices of time across the full exposure. It is clear that the amplitude of the fit is lower for later slices. It seems difficult, therefore, to attribute the DAMA signal to an induced modulation coming from exponentially decaying background activity.
@@
Figure 7: The result of performing the same fit as in Fig. 1 to a single realisation of a toy DAMA/LIBRA dataset, but this time including a component with a genuinely oscillating event rate, as described in the text. We compare a fit to only the modulating component of the datset (blue), versus the full dataset (orange), under two assumptions on the initial activities of 3H and 210Pb. Top: initial activities of 3H and 210Pb are set to those reported by DAMA. Bottom: initial activities of 3H and 210Pb are set to zero.